# Bayesian Multiobject Tracking With Neural-Enhanced Motion and Measurement Models

### v2

@@ Abstract
Multiobject tracking (MOT) is an important task in applications including autonomous driving, ocean sciences, and aerospace surveillance. Traditional MOT methods are model-based and combine sequential Bayesian estimation with data association and an object birth model. More recent methods are fully data-driven and rely on the training of neural networks. Both approaches offer distinct advantages in specific settings. In particular, model-based methods are generally applicable across a wide range of scenarios, whereas data-driven MOT achieves superior performance in scenarios where abundant labeled data for training is available. A natural thought is whether a general framework can integrate the two approaches. This paper introduces a hybrid method that utilizes neural networks to enhance specific aspects of the statistical model in Bayesian MOT that have been identified as overly simplistic. By doing so, the performance of the prediction and update steps of Bayesian MOT is improved. To ensure tractable computation, our framework uses belief propagation to avoid high-dimensional operations combined with sequential Monte Carlo methods to perform low-dimensional operations efficiently. The resulting method combines the flexibility and robustness of model-based approaches with the capability to learn complex information from data of neural networks. We evaluate the performance of the proposed method based on the nuScenes autonomous driving dataset and demonstrate that it has state-of-the-art performance.
Index Terms:
Multiobject tracking, Bayesian framework, neural networks, LiDAR, autonomous driving..
@@ I Introduction
  • Model-based methods face significant challenges using complex shape feature information of objects, e.g., as provided by LiDAR or camera measurements. In particular, it turns out to be challenging to describe this type of feature information using rather simple statistical models. In addition, for tractable modelling and estimation, model-based methods typically assume that objects move independently following Markovian state-transition models. This modelling approach ignores obvious complex interactions of object motion, as, e.g., induced by parallel lanes or intersections in an autonomous driving scenario.
@@
  • Data-driven MOT methods typically ignore expert knowledge or any physical description of the underlying processes representing object motion and sensing. While this “a priori information” effectively constrains the problem, it can also provide a high degree of interpretability. The absence of any model-based information leads to a lack of interpretability of data-driven MOT methods. Furthermore, most data-driven MOT methods cannot provide any uncertainty quantification of estimation results, which may be limiting in real-world applications.
@@
Figure 2: Block diagram of the proposed hybrid MOT approach that uses a neural network to enhance statistical motion and measurement models.
@@
In this paper, we address limitations of model-based and data-driven estimation by integrating neural networks into the statistical model used for Bayesian MOT. Using LiDAR and camera data from an autonomous driving scenario, we demonstrate that the resulting hybrid Bayesian method can improve the robustness and overall performance of MOT. The scenario used for performance evaluation, as well as a high-level block diagram of our approach, are shown in Figs. 1 and 2, respectively. The proposed method combines the strengths of well-designed and physically motivated statistical models with the ability of neural networks to extract complex feature information and complex interactions of object tracks. This combination is achieved by enhancing the established motion and measurement model of Bayesian MOT by the output of a neural network that has been trained using annotated data. Details on the neural-enhancement of motion and measurement model are provided in what follows.
@@
The neural-enhanced motion model is obtained by learning a state-transition function that captures the influence of historical object states and interactions with neighboring objects. The state transition function relies on a recurrent neural network (RNN) with hidden states and a weighting of neighbors inspired by the attention mechanism [11]. The neural-enhanced motion model is non-Markovian and makes it possible to consider temporal dependencies and spatial relationships between objects in the prediction step of MOT by propagating a high-dimensional hidden state that acts as a sufficient statistic for previous states of the object and objects in its vicinity. Sigma points (SPs) are used to perform the prediction step based on the neural-enhanced motion model efficiently [12, 13].
@@
A neural-enhanced measurement model is obtained by enhancing the model-based likelihood ratios used for data association and measurement update. Building on our previous work in [14, 15], our method provides two learnable factors referred to as “affinity factors” and “false positive rejection factors”. The affinity factors complement the affinity information provided by model-based likelihood ratios. In particular, affinity factors act as a surrogate for a statistical model of object size and shapes, which has been identified as notoriously challenging to establish. The false positive rejection factor dynamically adjusts the false positive PDF, individually for each measurement, using shape information extracted from raw sensor data. The neural network that provides these two factors can use different types of information as input, i.e., motion information provided by the prediction step, size measurements provided by the detector, and shape and appearance features learned from raw data.
@@
The overall structure of the statistical model is represented by a factor graph [16], where certain factors representing motion and measurement model are enhanced by the neural network. Our method follows the belief propagation framework for MOT [4] combined with sequential Monte Carlo (SMC) techniques [17] for a fast and accurate computation of marginal posterior PDFs. The number of objects and their states are finally computed based on these marginal posterior PDFs. The proposed method is applied to the nuScenes autonomous driving dataset [18] and can achieve state of-the-art object tracking performance.
@@ II Related Work and Contributions
Traditionally, model-based MOT methods can be further divided into two types. The first type completely relies on Bayesian estimation. Here, object appearance and departure, object motion, data association, and measurement generation are all described by well-established statistical models. This type of methods are referred to as Bayesian MOT methods. Representative methods include vector-type methods such as the joint probabilistic data association (JPDA) filter[3], the multiple hypothesis tracker (MHT)[19], and belief propagation (BP) [20, 4, 21, 22, 23], as well as set-type methods [5] such as probability hypothesis density (PHD) filters [6, 24, 25], the labeled multi-Bernoulli (LMB) filters [7, 26, 27], and Poisson multi-Bernoulli (PMB) filters [8, 4, 28].
To maintain computational tractability, model-based methods often adopt Gaussian mixture models or particle-based representations of PDFs [29, 6, 3]. In addition, to succeed in challenging real-world scenarios, Bayesian MOT methods that can adapt to changes in the statistical model by tracking additional nuisance parameters have been developed recently. In particular, methods that handle time-varying detection probabilities [30, 31], motion models [31, 32], and object extents [33, 34] have been introduced. Of particular interest for the proposed work is the BP approach for Bayesian MOT [20, 4, 21], which relies on a factor graph representation of the underlying statistical model. The factor graph representation facilitates the identification of parts of the statistical model that are overly simplistic and thus provide an opportunity for neural enhancement [14, 15].
The second type of model-based MOT methods combines sequential Bayesian estimation with deterministic techniques or heuristic rules for data association and track management [35, 36, 37, 38]. These methods typically use hand-designed rules or deterministic metrics, such as the Intersection-over-Union (IoU) and its variants (e.g., GIoU and CIoU), to compute association scores, representing the affinity of objects and measurements. Methods like the Hungarian algorithm or greedy matching are then commonly used to compute hard, i.e., non-probabilistic associations. Once associations are established, the Kalman filter[39] is typically adopted to perform sequential Bayesian estimation for the computation of marginal posterior PDFs of object states. This type of model-based MOT methods also perform track management based on heuristic rules.
@@
A straightforward strategy for combining model-based and data-driven approaches for MOT is to use data-driven processing and model-based estimation for different very specific tasks. For example, in [52], a neural network is used for track management and data association, while a Kalman filter is employed for sequential state estimation. Similarly, [53] uses feature measurements extracted by a neural network to leverage the complex position-related information of multipath components in an indoor localization scenario.
Our previously proposed neural-enhanced belief propagation (NEBP) method [14, 15] overcomes this limitation by tightly integrating data-driven and model-based processing for probabilistic data association. The main idea of NEBP for MOT is to introduce a GNN that has the same structure as the loopy subgraph used for probabilistic data association in BP-based MOT, and to exchange BP and GNN messages between the two graphs. Since the GNN has been trained using labeled data and makes use of shape features learned from raw data, NEBP is expected to lead to an improved data association solution and thus to an improved overall MOT performance [54]. However, the loopy structure of NEBP complicates the training of the GNN, and the message exchange between BP and GNN also leads to a lack of theoretical guarantees for the resulting loopy message passing algorithm. In addition, NEBP does not neural-enhance the prediction step of MOT. Contrary to existing data-driven approaches that only aim at improving object state prediction [46, 47, 48, 46, 49], we employ a learnable, non-Markovian motion model for calculating predicted PDFs within a Bayesian MOT framework.
The presented neural-enhanced Bayesian method for MOT advances over NEBP for MOT by directly enhancing the underlying motion and measurement models. It furthermore leads to a significantly improved MOT performance in real-world problems. The key contributions of this paper are summarized as follows.
@@ III Background on Model-Based MOT
In this section, we provide necessary background on model-based MOT. If not indicated otherwise, we follow the conventional statistical model for MOT presented in [4]. Before discussing the statistical model in detail, we discuss basic operations performed by model-based MOT.
@@
The prediction step computes predicted PDFs $f(\bm{\mathbf{y}}^{i}_{k}|$ $\bm{\mathbf{z}}_{1:k-1})$ based on marginal posterior PDFs $f(\bm{\mathbf{y}}^{i}_{k-1}|\bm{\mathbf{z}}_{1:k-1})$ for $i\in\{1,\dots,I_{k}\}$ from the previous time step by making use of the motion model. The measurement update step, based on Bayes rule, is used to then obtain the marginal posterior PDFs, i.e.,
@@
Here $l(\bm{\mathbf{z}}_{k}|\bm{\mathbf{y}}^{i}_{k})$ is a representation of the measurement model. Due to measurement-origin uncertainty, the functions $l(\bm{\mathbf{z}}_{k}|\bm{\mathbf{y}}^{i}_{k})$ are coupled across objects. (This coupling is currently not indicated by our simplified notation.) Finally, for $i\in\{I_{k}+1,\dots,N_{k}\}$ the object initialization step can be denoted as
@@
Finally, Bayesian state estimation for each object state $\bm{\mathbf{y}}^{i}_{k}$, $i\in\{1,\dots,N_{k}\}$ is performed by applying, e.g., the minimum mean square error estimator (MMSE). Bayesian estimation always relies on marginal PDFs $f(\bm{\mathbf{y}}^{i}_{k}|\bm{\mathbf{z}}_{1:k})$.
@@ III-A Object States and Motion Model
At time $k$, there are $N_{k}$ potential objects (POs) with kinematic states $\bm{\mathbf{x}}^{i}_{k}$, $i\in\{1,2,\dots,N_{k}\}$, where $N_{k}$ is the maximum possible number of objects that have generated a measurement up to time $k$ [4]. The kinematic state of an object $\bm{\mathbf{x}}^{i}_{k}\in\mathbb{R}^{d_{\mathrm{ki}}}$ consists of the object’s position and further motion-related parameters. For example, in a 2-D MOT scenario, the kinematic state can be defined as $\bm{\mathbf{x}}^{i}_{k}=\big{[}(\bm{\mathbf{p}}^{i}_{k})^{\mathrm{T}}\hskip 0.85358pt(\bm{\mathbf{v}}^{i}_{k})^{\mathrm{T}}\big{]}^{\mathrm{T}}$, where $\bm{\mathbf{p}}^{i}_{k}\in\mathbb{R}^{2}$ and $\bm{\mathbf{v}}^{i}_{k}\in\mathbb{R}^{2}$ are the 2-D position and velocity, respectively.
[…]
$f\big{(}\bm{\mathbf{x}}^{i}_{k},r^{i}_{k}\!=0\big{)}=p\big{(}r^{i}_{k}\!=0\big{)}\hskip 0.85358ptf_{\mathrm{d}}\big{(}\bm{\mathbf{x}}^{i}_{k}\big{)}$ (3)
where $p\big{(}r^{i}_{k}\!=0\big{)}\in[0,1]$ is the probability that the $i$th object does not exist at time $k$ and $f_{\mathrm{d}}\big{(}\bm{\mathbf{x}}^{i}_{k}\big{)}$ is an arbitrary “dummy PDF”. A measurement detector, applied to the raw sensor data, produces $J_{k}$ measurements denoted as $\bm{\mathbf{z}}_{k}=[({\bm{\mathbf{z}}^{1}_{k}})^{\mathrm{T}}~{}({\bm{\mathbf{z}}^{2}_{k}})^{\mathrm{T}}\cdots({\bm{\mathbf{z}}^{J_{k}}_{k}})^{\mathrm{T}}]^{\mathrm{T}}$.
Generally, there are two types of PO states:
@@
    Legacy POs denoted as $\bm{\mathbf{y}}^{i}_{k}=\underline{\bm{\mathbf{y}}}^{i}_{k}=[(\underline{\bm{\mathbf{x}}}^{i}_{k})^{\mathrm{T}}~{}\underline{r}^{i}_{k}]^{\mathrm{T}},i\in\{1,2,\dots,I_{k}\}$ represent objects that generated a measurement at a previous time step $k^{\prime}<k$. The joint state of legacy POs is defined as $\underline{\bm{\mathbf{y}}}_{k}=[(\underline{\bm{\mathbf{y}}}^{1}_{k})^{\mathrm{T}}\cdots(\underline{\bm{\mathbf{y}}}^{I_{k}}_{k})^{\mathrm{T}}]^{\mathrm{T}}$.
@@
    New POs denoted as $\bm{\mathbf{y}}^{i}_{k}=\overline{\bm{\mathbf{y}}}^{j}_{k}=[(\overline{\bm{\mathbf{x}}}^{j}_{k})^{\mathrm{T}}~{}\overline{r}^{j}_{k}]^{\mathrm{T}}$, $i=j+I_{k}$, $j\in\{1,2,\dots,J_{k}\}$ represent objects that generated a measurement at time step $k$ for the first time. Consequently, a new PO state $\overline{\bm{\mathbf{y}}}^{j}_{k}$ is introduced for each measurement $\bm{\mathbf{z}}^{j}_{k},j\in\{1,2,\dots,J_{k}\}$. The joint state of new POs is defined as $\overline{\bm{\mathbf{y}}}_{k}=[(\overline{\bm{\mathbf{y}}}^{1}_{k})^{\mathrm{T}}\cdots(\overline{\bm{\mathbf{y}}}^{J_{k}}_{k})^{\mathrm{T}}]^{\mathrm{T}}$.
When the measurements of the next time step are considered, new POs will become legacy POs. Thus, the total number of POs is $N_{k}=I_{k}+J_{k}$ and the number of legacy POs at time $k$ is $I_{k}=I_{k-1}+J_{k-1}=N_{k-1}$. The joint PO states $\bm{\mathbf{y}}_{k}$ at time $k$ is defined as $\bm{\mathbf{y}}_{k}=[(\underline{\bm{\mathbf{y}}}_{k})^{\mathrm{T}}~{}(\overline{\bm{\mathbf{y}}}_{k})^{\mathrm{T}}]^{\mathrm{T}}$.
@@
If the PO $i$ does not exist at time $k-1$, i.e., $r_{k-1}^{i}=0$, then it cannot exist at time $k$ either. Thus, the state transition PDF can be expressed as
@@
If the PO $i$ exists at time $k-1$, i.e. $r_{k-1}^{i}=1$, then it continues to exist with a survival probability $p_{s}$, i.e.,
@@
=$\hskip 22.76219pt=\begin{cases}p_{s}f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k_{i}:k-1},{\mathbf{y}}^{\sim i}_{k_{i}:k-1}),&\hskip 4.2679pt\underline{r}^{i}_{k}=1\\[2.84526pt] (1-p_{s})f_{\mathrm{d}}(\underline{\bm{\mathbf{x}}}^{i}_{k}),&\hskip 4.2679pt\underline{r}^{i}_{k}=0.\end{cases}$ (8)
However, due to high-order dependencies, establishing a statistical model for the state-transition PDF $f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k_{i}:k-1},{\mathbf{y}}^{\sim i}_{k_{i}:k-1})$ in (8) is quite challenging. In the next section, we will elaborate on this in more detail.

### v3

@@ Abstract
Multiobject tracking (MOT) is an important task in applications including autonomous driving, ocean sciences, and aerospace surveillance. Traditional MOT methods are model-based and combine sequential Bayesian estimation with data association and an object birth model. More recent methods are fully data-driven and rely on the training of neural networks. Both approaches offer distinct advantages in specific settings. In particular, model-based methods are generally applicable across a wide range of scenarios, whereas data-driven MOT achieves superior performance in scenarios where abundant labeled data for training is available. A natural thought is whether a general framework can integrate the two approaches. This paper introduces a hybrid method that utilizes neural networks to enhance aspects of the statistical motion and measurement models in Bayesian MOT that have been identified as overly simplistic. To ensure tractable computation, our framework uses belief propagation (BP) to avoid high-dimensional operations, combined with sequential Monte Carlo and sigma point (SP) techniques to perform low-dimensional operations efficiently. The resulting MOT method combines the flexibility and robustness of model-based approaches with the capability to learn complex information from data of neural networks. We evaluate the performance of the proposed method based on the nuScenes autonomous driving dataset and demonstrate that it has state-of-the-art performance.
@@ I Introduction
  • Model-based methods face significant challenges using complex shape feature information of objects, e.g., as provided by LiDAR or camera measurements. In particular, it turns out to be challenging to describe this type of feature information using rather simple statistical models. Moreover, to maintain tractable modeling and estimation, model-based approaches typically assume that objects move independently according to Markovian state-transition models. As a result, explicit interactions in object motion–such as those induced by parallel lanes or intersections in autonomous driving scenarios–are commonly ignored.
@@
  • Data-driven MOT methods typically ignore expert knowledge or any physical description of the underlying processes representing object motion and sensing. While this “a priori information” effectively constrains the problem, it can also provide a high degree of interpretability. Its absence leads to data-driven MOT methods being less interpretable. Furthermore, most data-driven MOT methods cannot provide any uncertainty quantification of estimation results, which may be limiting in real-world applications.
@@
Figure 2: Block diagram of a single time step of the proposed MOT method that uses a neural network to enhance statistical motion and measurement models.
@@
In this paper, we address limitations of model-based and data-driven estimation by integrating neural networks into the statistical model used for Bayesian MOT. Using LiDAR and camera data from an autonomous driving scenario, we demonstrate that the resulting hybrid Bayesian method can improve the robustness and overall performance of MOT. The scenario used for performance evaluation, as well as a high-level block diagram of our approach, are shown in Figs. 1 and 2, respectively. The proposed method combines the strengths of well-designed and physically motivated statistical models with the ability of neural networks to extract complex feature information and complex interactions of object tracks. This combination is achieved by enhancing the established motion and measurement model of Bayesian MOT by the output of a neural network that has been trained using annotated data. Details on the neural-enhancement of motion and measurement model are provided in what
follows.
@@
The neural-enhanced motion model is obtained by learning a state-transition function that captures the influence of historical object states and interactions with neighboring objects. The state transition function relies on a recurrent neural network (RNN) with hidden states and a weighting of neighbors inspired by the attention mechanism [52]. The neural-enhanced motion model is non-Markovian and makes it possible to consider temporal dependencies and spatial relationships between objects in the prediction step of MOT by propagating a high-dimensional hidden state that acts as a sufficient statistic for previous states of the modeled object and objects in its vicinity. Sigma points (SPs) are used to perform the prediction step based on the neural-enhanced motion model efficiently [20, 37]
@@
A neural-enhanced measurement model is obtained by enhancing the model-based likelihood ratios used for data association and measurement update. Building on our previous work in [28, 63], our method provides two learnable factors referred to as “affinity factors” and “false positive rejection factors”. The affinity factors complement the affinity information provided by model-based likelihood ratios. In particular, affinity factors act as a surrogate for a statistical model of object size and shapes, which has been identified as notoriously challenging to establish. The false positive rejection factor dynamically adjusts the false positive probability density function (PDF), individually for each measurement, using shape information extracted from raw sensor data. The neural network that provides these two factors can use different types of information as input, i.e., motion information provided by the prediction step, size measurements provided by the detector, and shape and appearance features learned from raw data
@@
The overall structure of the statistical model is represented by a factor graph [23], where certain factors representing motion and measurement model are enhanced by the neural network. Our method follows the belief propagation (BP) framework for MOT [38] combined with sequential Monte Carlo and sigma point (SP) techniques [1] for a fast and accurate computation of marginal posterior PDFs. The number of objects and their states are finally computed based on these marginal posterior PDFs. We evaluate the proposed method based on LiDAR and camera data from the nuScenes autonomous driving dataset [10] and further assess its performance across various detectors and sensor modalities. In all tested configurations, our approach consistently achieves state-of-the-art performance.
@@ II Related Work and Contributions
Traditionally, model-based MOT methods can be further divided into two types. The first type completely relies on Bayesian estimation. Here, object appearance and departure, object motion, data association, and measurement generation are all described by well-established statistical models. This type of methods are referred to as Bayesian MOT methods. Representative methods include vector-type methods such as the joint probabilistic data association (JPDA) filter[3], the multiple hypothesis tracker (MHT)[7], and BP [35, 38, 34, 70, 53, 60], as well as set-type methods [32] such as probability hypothesis density (PHD) filters [55, 54], the labeled multi-Bernoulli (LMB) filters [45, 56, 42], and Poisson multi-Bernoulli (PMB) filters [67, 38, 14].
To maintain computational tractability, model-based methods often adopt Gaussian mixture models or particle-based representations of PDFs [21, 55, 3]. In addition, to succeed in challenging real-world scenarios, Bayesian MOT methods that can adapt to changes in the statistical model by tracking additional nuisance parameters have been developed recently. In particular, methods that handle time-varying detection probabilities [61, 49], motion models [49, 64], and object extents [62, 15] have been introduced. Of particular interest for the proposed work is the BP approach for Bayesian MOT [35, 38, 34], which relies on a factor graph representation of the underlying statistical model. The factor graph representation facilitates the identification of parts of the statistical model that are overly simplistic and provide an opportunity for neural enhancement [28, 63].
The second type of model-based MOT methods combines sequential Bayesian estimation with deterministic techniques or heuristic rules for data association and track management [43, 27, 59, 26]. These methods typically use hand-designed rules or deterministic metrics, such as the Intersection-over-Union (IoU) and its variants, to compute association scores, representing the affinity of objects and measurements. Methods like the Hungarian algorithm or greedy matching are then commonly used to compute hard, i.e., non-probabilistic associations. Once associations are established, the Kalman filter[18] is typically adopted to perform sequential Bayesian estimation for the computation of marginal posterior PDFs of object states. This type of model-based MOT methods also perform track management based on heuristic rules.
@@
A straightforward strategy for combining model-based and data-driven approaches for MOT is to use data-driven processing and model-based estimation for different very specific tasks. For example, in [12], a neural network is used for track management and data association, while a Kalman filter is employed for sequential state estimation. Our previously proposed neural-enhanced BP (NEBP) method [28, 63] directly integrates data-driven and model-based processing for probabilistic data association. The main idea of NEBP for MOT is to introduce a GNN that has the same structure as the loopy subgraph used for probabilistic data association in BP-based MOT, and to exchange BP and GNN messages between the two graphs. Since the GNN has been trained using labeled data and makes use of shape features learned from raw data, NEBP is expected to lead to an improved data association solution and thus to an improved overall MOT performance [47]. However, the loopy structure of NEBP complicates the training of the GNN, and the message exchange between BP and GNN also leads to a lack of theoretical guarantees for the resulting loopy message passing algorithm. In addition, NEBP does not neural-enhance the prediction step of MOT. Contrary to existing data-driven approaches that only aim at improving object state prediction [24, 44, 30, 24, 16], we employ a learnable, non-Markovian motion model with object interactions for calculating predicted PDFs within a Bayesian MOT framework.
The presented neural-enhanced Bayesian method for MOT advances over NEBP for MOT by directly enhancing the underlying motion and measurement models. The outputs of the neural network and the statistical models are combined in a unified framework to improve the computation of posterior PDFs and object state estimates. It furthermore leads to a significantly improved MOT performance in real-world problems. The key contributions of this paper are summarized as follows.
@@ III Background on Model-Based MOT
In this section, we provide necessary background on model-based MOT. If not indicated otherwise, we follow the conventional statistical model for MOT presented in [38]. Before discussing the statistical model in details, we introduce basic operations performed by model-based MOT.
@@
The prediction step computes predicted marginal PDFs $f(\bm{\mathbf{y}}^{i}_{k}|$ $\bm{\mathbf{z}}_{1:k-1})$ based on marginal posterior PDFs $f(\bm{\mathbf{y}}^{i}_{k-1}|\bm{\mathbf{z}}_{1:k-1})$ for $i\in\{1,\dots,I_{k}\}$ from the previous time step by making use of the motion model. The measurement update step, based on Bayes rule, is used to then obtain the marginal posterior PDFs, i.e.,
@@
Here $l(\bm{\mathbf{z}}_{k}|\bm{\mathbf{y}}^{i}_{k})$ is a statistical representation of the measurement model. Due to measurement-origin uncertainty, the functions $l(\bm{\mathbf{z}}_{k}|\bm{\mathbf{y}}^{i}_{k})$ are coupled across objects. (This coupling is currently not indicated by our simplified notation.) Finally, for $i\in\{I_{k}+1,\dots,N_{k}\}$ the object initialization step can be denoted
as
@@
Finally, Bayesian state estimation for each object state $\bm{\mathbf{y}}^{i}_{k}$, $i\in\{1,\dots,N_{k}\}$ is performed by applying, e.g., the minimum-mean-square-error (MMSE). Bayesian estimation always relies on marginal PDFs $f(\bm{\mathbf{y}}^{i}_{k}|\bm{\mathbf{z}}_{1:k})$.
@@ III-A Object States and Motion Model
At time $k$, there are $N_{k}$ potential objects (POs) with kinematic states $\bm{\mathbf{x}}^{i}_{k}$, $i\in\{1,2,\dots,N_{k}\}$, where $N_{k}$ is the maximum possible number of objects that have generated a measurement up to time $k$ [38]. The kinematic state of an object $\bm{\mathbf{x}}^{i}_{k}\in\mathbb{R}^{d_{\mathrm{ki}}}$ consists of the object’s position and further motion-related parameters. For example, in a 2-D MOT scenario, the kinematic state can be defined as $\bm{\mathbf{x}}^{i}_{k}=\big[(\bm{\mathbf{p}}^{i}_{k})^{\mathrm{T}}\hskip 0.85358pt(\bm{\mathbf{v}}^{i}_{k})^{\mathrm{T}}\big]^{\mathrm{T}}\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt\hskip-0.85358pt$, where $\bm{\mathbf{p}}^{i}_{k}\in\mathbb{R}^{2}$ and $\bm{\mathbf{v}}^{i}_{k}\in\mathbb{R}^{2}$ are the 2-D position and velocity, respectively.
The existence of each PO state is modeled by a binary random variable $r^{i}_{k}\in\{0,1\}$. A PO state $\bm{\mathbf{x}}^{i}_{k}$ exists if and only if $r^{i}_{k}=1$. For simplicity, we use the notation ${\bm{\mathbf{y}}}^{i}_{k}=[(\bm{\mathbf{x}}^{i}_{k})^{\mathrm{T}}\hskip 0.85358ptr^{i}_{k}]^{\mathrm{T}}$ to indicate a PO state that has been augmented by an existence variable. The kinematic state $\bm{\mathbf{x}}^{i}_{k}$ of a nonexisting PO is obviously irrelevant. More specifically, for $r^{i}_{k}=0$, all PDFs defined for augmented PO states, $f\big(\bm{\mathbf{y}}^{i}_{k}\big)=f\big(\bm{\mathbf{x}}^{i}_{k}\hskip-0.85358pt,r^{i}_{k}\big)$, can be expressed as $f\big(\bm{\mathbf{x}}^{i}_{k}\hskip-0.85358pt,r^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt0\big)=p\big(r^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt0\big)\hskip 0.85358ptf_{\mathrm{d}}\big(\bm{\mathbf{x}}^{i}_{k}\big)$ where $p\big(r^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt0\big)\hskip-0.85358pt\in\hskip-0.85358pt[0,1]$ is the probability that the $i$th object does not exist at time $k$ and $f_{\mathrm{d}}\big(\bm{\mathbf{x}}^{i}_{k}\big)$ is an arbitrary “dummy PDF”. Similarly, for $r^{i}_{k}=1$, PDFs defined for augmented PO states are given by $f\big(\bm{\mathbf{x}}^{i}_{k}\hskip-0.85358pt,r^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt1\big)=p\big(r^{i}_{k}\!\hskip-0.85358pt=\hskip-0.85358pt1\big)\hskip 0.85358ptf\big(\bm{\mathbf{x}}^{i}_{k}\big)$, where $f(\bm{\mathbf{x}}^{i}_{k})$ is the PDF of kinematic state $\bm{\mathbf{x}}^{i}_{k}$. A measurement detector, applied to the raw sensor data, produces $J_{k}$ measurements denoted as $\bm{\mathbf{z}}_{k}=[({\bm{\mathbf{z}}^{1}_{k}})^{\mathrm{T}}~({\bm{\mathbf{z}}^{2}_{k}})^{\mathrm{T}}\cdots({\bm{\mathbf{z}}^{J_{k}}_{k}})^{\mathrm{T}}]^{\mathrm{T}}\hskip-0.85358pt$
Generally, there are two types of PO states
@@
    Legacy POs denoted as $\bm{\mathbf{y}}^{i}_{k}=\underline{\bm{\mathbf{y}}}^{i}_{k}=[(\underline{\bm{\mathbf{x}}}^{i}_{k})^{\mathrm{T}}~\underline{r}^{i}_{k}]^{\mathrm{T}},i\in\{1,2,\dots,I_{k}\}$ represent objects that generated a measurement at a previous time step $k^{\prime}<k$. The joint state of legacy POs is defined as $\underline{\bm{\mathbf{y}}}_{k}=[(\underline{\bm{\mathbf{y}}}^{1}_{k})^{\mathrm{T}}\cdots(\underline{\bm{\mathbf{y}}}^{I_{k}}_{k})^{\mathrm{T}}]^{\mathrm{T}}\hskip-0.85358pt\hskip-0.85358pt$
@@
    New POs denoted as $\bm{\mathbf{y}}^{i}_{k}=\overline{\bm{\mathbf{y}}}^{j}_{k}=[(\overline{\bm{\mathbf{x}}}^{j}_{k})^{\mathrm{T}}~\overline{r}^{j}_{k}]^{\mathrm{T}}\hskip-0.85358pt\hskip-0.85358pt$, $i\hskip-0.85358pt=\hskip-0.85358ptj\hskip-0.85358pt+\hskip-0.85358ptI_{k}$, $j\in\{1,2,\dots,J_{k}\}$ represent objects that generated a measurement at time step $k$ for the first time. Consequently, a new PO state $\overline{\bm{\mathbf{y}}}^{j}_{k}$ is introduced for each measurement $\bm{\mathbf{z}}^{j}_{k},j\in\{1,2,\dots,J_{k}\}$. The joint state of new POs is defined as $\overline{\bm{\mathbf{y}}}_{k}=[(\overline{\bm{\mathbf{y}}}^{1}_{k})^{\mathrm{T}}\cdots(\overline{\bm{\mathbf{y}}}^{J_{k}}_{k})^{\mathrm{T}}]^{\mathrm{T}}\hskip-0.85358pt\hskip-0.85358pt$.
When the measurements of the next time step are considered, new POs will become legacy POs. Thus, the total number of POs is $N_{k}=I_{k}+J_{k}$ and the number of legacy POs at time $k$ is $I_{k}=I_{k-1}+J_{k-1}=N_{k-1}$. The joint PO state $\bm{\mathbf{y}}_{k}$ at time $k$ is defined as $\bm{\mathbf{y}}_{k}=[(\underline{\bm{\mathbf{y}}}_{k})^{\mathrm{T}}~(\overline{\bm{\mathbf{y}}}_{k})^{\mathrm{T}}]^{\mathrm{T}}$.
@@
If the PO $i$ does not exist at time k1k-1, i.e., $r_{k-1}^{i}=0$, then it cannot exist at time $k$ either. Thus, the state transition PDF can be expressed
as
@@
If the PO $i$ exists at time k1k-1, i.e. $r_{k-1}^{i}=1$, then it continues to exist with a survival probability $p_{s}$, i.e.
@@
$\hskip 28.45274pt=\begin{cases}p_{s}f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k_{i}:k-1},{\mathbf{y}}^{\sim i}_{k_{i}:k-1}),&\hskip 4.2679pt\underline{r}^{i}_{k}=1\\[2.84526pt] (1-p_{s})f_{\mathrm{d}}(\underline{\bm{\mathbf{x}}}^{i}_{k}),&\hskip 4.2679pt\underline{r}^{i}_{k}=0.\end{cases}$ (7)
However, due to high-order dependencies, establishing a statistical model for the state-transition PDF $f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k_{i}:k-1},{\mathbf{y}}^{\sim i}_{k_{i}:k-1})$ in (7) is quite challenging. This complicates the computation of the predicted PDF. For this reason, most existing MOT methods assume that state transition is performed based on a Markovian state transition model and the state transition of object $i$ is independent of the state transition of all the other objects, i.e., the model in (7) is simplified according to [4, 6, 3, 32, 38, 55, 45, 67]
$f(\underline{\bm{\mathbf{x}}}^{i}_{k},\underline{r}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k-1},1)=\begin{cases}p_{s}f({\underline{\bm{\mathbf{x}}}}^{i}_{k}|{\bm{\mathbf{x}}}^{i}_{k-1}),&\hskip 4.2679pt\underline{r}^{i}_{k}=1\\[2.84526pt] (1-p_{s})f_{\mathrm{d}}(\underline{\bm{\mathbf{x}}}^{i}_{k}),&\hskip 4.2679pt\underline{r}^{i}_{k}=0.\end{cases}$
A major goal of our proposed MOT approach is to avoid the Markovian state-transition model and the assumption that objects move independently